OpenStudy (anonymous):

4 2/5 + 3 3/5 ?? simplified

OpenStudy (anonymous):

\[4\frac{2}{5}+3\frac{3}{5}\]

OpenStudy (anonymous):

is that the question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so:\[4+3+\frac{2+3}{5}=7+\frac{5}{5}=7+1=8\]

OpenStudy (anonymous):

thats what i got too but wasnt sure. much appericated

OpenStudy (anonymous):

no prob